QUESTION 20

  1. You want to buy a new pickup for $43,696, and the finance office at the dealership has quoted you an 8.4% APR loan for 6 years of monthly payments. What is the effective annual rate on this loan? (Do not include the percent sign (%). Enter rounded answer as directed, but do not use the rounded numbers in intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)
Answer #1

Effective annual rate is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.0840/12)^12-1

        = 1.0873 -1

        = 0.0873*100

        = 8.7311%   8.73.

In case of any query, kindly comment on the solution.